OpenStudy (anonymous):

The function f(x) = 2^x is a logarithmic function. True or false

OpenStudy (whpalmer4):

Does it look like a logarithmic function to you?

OpenStudy (anonymous):

True ?

OpenStudy (whpalmer4):

What does a logarithmic function have in it?

OpenStudy (mathstudent55):

Exponential function: \(f(x) = k^x\) Logarithmic function: \(f(x) = \log_b x\)

OpenStudy (anonymous):

I think it's false am I right ?

OpenStudy (rock_mit182):



OpenStudy (whpalmer4):

A logarithmic function will have a logarithm in it :-)

OpenStudy (anonymous):

So it's true ?

OpenStudy (whpalmer4):

No, it is not true. You have an exponential function there. A logarithmic function will have \(\log\) or \(\ln\) in it somewhere. Here's a plot of what you have, compare it with the drawings @rock_mit18 provided:

OpenStudy (mathstudent55):

In the general exponential and logarithmic equations I wrote above, notice the exponential equation is just like yours when you replace k with 2.

OpenStudy (anonymous):

Can you still help me ?